Read and understand the problem below calculate the cost of goods sold and gross profit.

Shoe keeper had bought 100 footballs for ph.30.00 each.One football player comes and bouhgt 10 footballs for 50.00 each. Calculate the cost of goods sold. Show your solution
1
Expert's answer
2021-09-28T01:18:35-0400

cost of 10 balls ="10\times 30=300"

money received for selling 10 balls"=10\times 50=500"

gross profit =500-300= ph. 200